OpenStudy (anonymous):

find the solution to: 64^(4-x) = 4^(2x) & use complete sentences on how to solve it

jimthompson5910 (jim_thompson5910):

Hint: Rewrite 64 as 4^3 to get 64^(4-x) = 4^(2x) (4^3)^(4-x) = 4^(2x) Then multiply the exponents 4^(3(4-x)) = 4^(2x) Since the bases are equal, the exponents are equal, so... 3(4-x) = 2x

OpenStudy (anonymous):

thanks so much (:

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

tell me what you get for x

OpenStudy (anonymous):

3(4-x) = 12 - 3x (i FOILED on that one) then i got 12x - 3x = 2x . subtracted 12x & 3x to get 9x = 2x . then i divided by 2x & got 9/2 = 0 ??? that doesnt work :/ what did i do wrong ?

jimthompson5910 (jim_thompson5910):

it should be 12 - 3x and NOT 12x - 3x

jimthompson5910 (jim_thompson5910):

So it should be 12 - 3x = 2x Try it again

OpenStudy (anonymous):

oh i added an extra x . so its 12 = 5x . and then 12/5 = x ?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

great ! :D thanks!

jimthompson5910 (jim_thompson5910):

sure thing